# BAP v2: An Enhanced Task Framework for Instruction Following in Minecraft Dialogues

### v2

@@ 3.2 Creating a clean test set
The BAP v2 test set consists of 1,155 BAP items, all of which have a clear context and a correct built structure. 1,071 of these items have a unique correct interpretation, and 84 of the empty board (EB) items have multiple correct interpretations. This set was obtained from the original BAP test set of 1,616 BAP items by removing 461 items that should not be used for evaluation (418 with unclear contexts and 43 with incorrect structures that could not be fixed), and by fixing the structures of 149 items whose original structures were incorrect.
@@ 4.1 Fairer F1
The clean v2 test set consists only of BAP items with a clear context and correct structures, but 84 of the EB items (7.3% of the v2 test set) do not have a unique interpretation, since their structure can be placed and oriented in many different ways (see Figure 7 above). The v2 test set can therefore be split up into items with a unique correct interpretation ($\mathcal{U}$) and items with multiple correct interpretations ($\mathcal{M}$). But while the strict F1 score defined in Section 2.4 is appropriate for items in $\mathcal{U}$, it unfairly penalizes Builders on items in $\mathcal{M}$ when their structure is correct, but placed in a different location and/or orientation than the reference structure. A fair evaluation metric needs to account for these allowable changes in location and orientation for items in $\mathcal{M}$, i.e. structures placed on an empty board when the context does not uniquely identify their placement.
@@
To evaluate a model’s net action set $A_{M}^{\mathrm{net}}$ against a human reference net action set $A_{H}^{\mathrm{net}}$ for an item in $\mathcal{M}$, where $A_{M}$ leads from the (empty) board to a predicted structure $S_{M}$, and $A_{H}$ leads from the (empty) board to a reference structure $S_{H}$, we therefore first identify an optimal alignment of the predicted structure to the reference, $\mathcal{A}^{*}=\operatorname*{arg\,min}_{\mathcal{A}}(\Delta(\mathcal{A}(S_{M}),S_{H}))$, and apply the same transformation $\mathcal{A}^{*}$ to all actions in the model’s net action set $A_{M}^{\mathrm{net}}$, yielding an aligned action sequence $\widetilde{A}_{M}^{\mathrm{net}}$ that can now be fairly compared against $A_{H}^{\mathrm{net}}$ by our original strict F1 metric. That is, a fairer F1 metric computes a strict F1 score for the predicted action sequences $A_{M}^{\mathrm{net}}$ for any item in $\mathcal{U}$, because the reference sequences for these items yield the only correct structures for these items, but for any item in $\mathcal{M}$, it computes a strict F1 score for the optimally aligned predicted action sequences $\widetilde{A}_{M}^{\mathrm{net}}$ because this item has multiple correct interpretations that only differ from each other in location and orientation. Thus, in the example in Figure 7, any predicted orange 3x3 U-shaped structure placed on the ground will receive a fairer F1 score of 1.0.
@@ 4.2 Auxiliary metrics
[…]
@@
$M^{\textrm{Type}}$ =$=\{t\mid(t,c,\langle x,y,z\rangle)\in A^{\textrm{net}}\}$
$M^{\textrm{Color}}$ =$=\{(t,c)\mid(t,c,\langle x,y,z\rangle)\in A^{\textrm{net}}\}$
$M^{\textrm{Loc}}$ =$=\{\langle x,y,z\rangle\mid(t,c,\langle x,y,z\rangle)\in A^{\textrm{net}}\}$
which yields Type, Color and Location (strict) precision, recall and F1 scores by comparing the sizes of the corresponding multisets and their intersections (we only show Type F1 below, as Color and Location F1 are similarly defined):
Strict Type $\textbf{Strict Type Precision}~{}P_{\textrm{Type}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{|(M_{M}^{\textrm{Type}}\cap M_{H}^{\textrm{Type}})|}{|M_{M}^{\textrm{Type}}|}$
Strict Type $\textbf{Strict Type Recall}~{}R_{\textrm{Type}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{|(M_{M}^{\textrm{Type}}\cap M_{H}^{\textrm{Type}})|}{|M_{H}^{\textrm{Type}}|}$
[…]
Intuitively, the Type F1 metric quantifies how good a model is at understanding what type of actions to perform – placements, or removals, or both, and how many of each type. The Color F1 metric narrows down Type F1 a bit more to factor for colors as well. It quantifies how good a model is at understanding the colors of blocks that need to be placed or removed. The Location F1 metric quantifies how good a model is at understanding the specific locations of blocks that need to be placed or removed (without regard to color), i.e., spatial reasoning. Also, as mentioned earlier, these metrics are analogous to the strict F1 metric. They can then be used to define their fairer counterparts in the same way the strict F1 was used to define the fairer F1. Henceforth, we will use Type, Color and Location F1 to refer to their fairer counterparts.
@@
To evaluate how good a Builder model is at understanding the "shape" in which the required blocks need to be placed or removed, we define a metric, Shape F1, that is agnostic to the absolute $\langle x,y,z\rangle$ locations of the required blocks, but only cares that they be placed/removed in the right locations relative to each other. Essentially, it compares net actions modulo their exact placement and orientation. Therefore, similar to the methodology used for fairer F1 (Section 4.1), we optimally align the model’s net actions $A_{M}^{\mathrm{net}}$ against the human reference net actions $A_{H}^{\mathrm{net}}$ (instead of the corresponding resulting built structures themselves like was done for fairer F1), yielding aligned net actions $\hat{A}_{M}^{\mathrm{net}}$ that can now be compared against $A_{H}^{\mathrm{net}}$ by our original strict F1 metric. This is done for all items in the v2 test set (and not just the ones with multiple interpretations like was done for fairer F1). We will see examples of Shape F1 and the other auxiliary metrics later in Sections 7.2 and 8.1.
4.3 Micro vs Macro F1
Current BAP evaluation reports the micro-averaged F1 (micro F1) over all action sequences in the test data (Section 2.4), reflecting a model’s quality in predicting individual actions. However, since the BAP task requires sequence prediction, it’s equally important to assess the model’s quality in predicting entire action sequences using the macro-averaged F1 (macro F1). Intuitively, this is also particularly relevant for Shape F1, which measures the model’s performance on a broader sequence-level property (the "shape" of the net actions). Therefore, hereafter, we will report both micro and macro F1.
@@ 4.4 BAP v2 evaluation benchmark
The above metrics enable a fairer, robust, and more insightful evaluation on the clean v2 test set. We use the fairer F1 metric to report both micro and macro F1 scores in Tables 1 and 2 respectively. For each, we include both auxiliary and overall F1 metrics, and report them on the overall/full test set and separately for the EB and NEB subsets as well. This forms the BAP v2 evaluation benchmark. Our approach also enables more accurate evaluation within the EB and NEB categories, offering a clearer understanding of model performance across these two distinct scenarios. This division is important for the following reasons:
  • EB examples are generally the easiest for the BAP task, making EB performance a basic competency or sanity test, and it also reflects a model’s maximum potential quality.
  • EB performance is also an estimate of the ability of a model to begin a game accurately – an elementary yet important trait to measure. Moreover, in an interactive/online setting, from a user (human Architect) experience perspective and considering the potential for subsequent cascading errors, mistakes made in EB scenarios can potentially be costlier than those in NEB, as EB marks the start of a game (even though EB examples in the data are far fewer than NEB). Quantifying EB performance will help provide a sense for how a model can potentially fare in an interactive setting wrt this aspect.
We begin with the micro scores in Table 1 for direct comparison with our prior work. Recall that the baseline model achieves an F1 score of 21.1% under the strict F1 metric. This increases to 27.3% under the fairer F1 metric (the overall F1 on the overall dataset), a notable difference of 6.2 percentage points and a 29.4% improvement. This highlights the importance of the high quality cleaner test data and the fairer F1 metric in providing a fairer reflection of model capability. On the overall dataset, the model achieves reasonably good F1 scores of 65.2% for Type and 62.4% for Color, but only 27.7% for Location F1, which is also very close to the overall F1 of 27.3%. This suggests that spatial reasoning (Location F1) is the key bottleneck for high performance on this task. As expected, the model performs better on EB than NEB across all metrics. Recall that the number of EB items are fewer than NEB items though (and the metrics on the overall dataset will therefore be more influenced by NEB). Next, we turn to the macro scores in Table 2. While all values increase compared to the micro scores, the overall trends remain the same. The overall F1 on the overall test set increases from 27.3% (micro) to 28.1% (macro). The model achieves a Shape F1 of 36.3% (micro) and 41.1% (macro) on the overall test set, with the macro score providing a more intuitive measure in this context, as discussed in Section 4.3.
Hereafter, we will use the v2 evaluation to report model performance.
@@
Table 1: Micro F1 scores for the baseline model under the v2 benchmark
Dataset Type Color Location Shape Overall
EB 80.9 79.8 61.5 64.9 61.2
NEB 73.1 69.9 25.7 39.0 25.2
Overall 73.8 70.7 28.5 41.1 28.1
Table 2: Macro F1 scores for the baseline model under the v2 benchmark
5 Revisiting BAP Data: Generating Synthetic Data
One key challenge of the BAP task is the limited training data, despite the data augmentation (Section 2.5). We aim to avoid simply collecting more data like the MDC, due to scalability issues and the desire to stay aligned with many realistic scenarios, where data collection for complex situated dialogue tasks, like the MCBT, can be impractical/expensive, warranting exploration of alternative methods. Additionally, we observe that the baseline model struggles primarily with spatial reasoning, as reflected in its Location F1 score. These reasons motivate the need for additional synthetic training data that, while naturally simpler than the MDC, remains rich in spatial relations and referential expressions to help train more robust models. Looking ahead, such data could also be crucial for training more sophisticated, data-hungry deep transformer models and training/fine-tuning increasingly large LLMs. (We, in fact, also illustrate the impact of this data on training a simple LLM and transformer-based model in Section 6.3.)
@@ 5 Revisiting BAP Data: Generating Synthetic Data
We design novel Minecraft dialogue and target structure simulators that emulate the MCBT, and use them to generate a set of three synthetic dialog datasets with corresponding target structures . Each dataset is modeled after the MDC and is understandably simpler, while preserving essential elements of realistic human behavior observed in the MDC as much as possible, such as the language, clarification exchanges, Architect A ’s and Builder B ’s planning perspectives, dialog flow, etc. Across these datasets, we vary the level of instruction abstraction, the complexity of spatial relations and referential expressions, the complexity of target structures, etc. Special emphasis is laid on spatial relations as that is a core part of our motivation here. Each dataset is thus unique, contributing to the overall diversity and richness of the overall set.
@@
Note that although our primary focus is on BAP, the simulators and data introduced here are applicable to other MCBT subtasks as well (e.g. the AUG subtask; Section 2.1.2), supporting advancements across the broader MCBT ecosystem. We discuss this in more detail in Section 8.4.
5.1 General dialog simulation framework
5.1.1 Desiderata
Similar to the target structures and game logs in the MDC, the following components of synthetic dialogues need to be generated:
  • The target structure
  • The dialogue/game – a chronological sequence that interleaves the following elements:
    • A dialogue between A and B, consisting of utterances that include instructions, potential clarification exchanges, etc.
    • B’s position and orientation at appropriate points in the game, i.e., when an utterance or action is recorded (Grounded elements of the dialog, e.g. spatial relations, are always wrt this frame of reference.)
    • B’s actions, which reflect the evolving structure being built
5.1.2 High-level Algorithm
As described in Section 2.1, the MCBT task is asynchronous and loosely structured. The MDC dialogs are real human-human game logs collected in a low-stakes, free-flowing setting with minimal constraints on participants’ language and actions. To enable tractable simulation and facilitate simpler synthetic data generation, we make some simplifications and structure the task as follows.
Our dialog simulators generate synthetic dialogs that meet the desiderata, following a general framework. At a high level, it is an iterative procedure that continues until the target structure is built. The target structure may be provided as input or generated dynamically. In the former case, a separate simulator generates target structures. Each iteration produces one dialog segment—A’s instruction followed by B’s actions in response, with optional clarification exchanges and confirmations—leading to the construction of a part of the target structure. We are essentially simulating A and B, including their latent planning processes. Each iteration proceeds as shown below. (We use Figure 11 as an example. It illustrates a BAP item $(\mathcal{H}^{t-1},S^{t-1},D^{t},A^{t},S^{t})$ from one of our synthetic datasets (recall notation from Section 2.2). In this iteration, the dialog and actions following $S^{t-1}$ are simulated to produce the next structure, $S^{t}$ – specifically, simulating $D^{t}$, $A^{t}$, and $S^{t}$.)

### v3

@@ 3.2 Statistics of the clean BAP v2 test set
To obtain our final evaluation set, we start with the original BAP test set of 1,616 BAP items, but correct the structures of 149 items that we were able to identify and fix automatically, and remove 461 items that are unsuitable for evaluation purposes (418 with unclear contexts and 43 with incorrect structures that could not be fixed). This results in a new, BAP v2 test set that consists of 1,155 BAP items, all of which have a clear context and a correct built structure. 1,071 of these items have a unique correct interpretation, and 84 of the empty board (EB) items have multiple correct interpretations.
@@ 4.1 Fairer F1
The clean v2 test set consists only of BAP items with a clear context and correct structures. 84 of the EB items (7.3% of the v2 test set) do not have a unique interpretation, since their structure can be placed and oriented in many different ways (Section 3.1). The v2 test set can therefore be split up into items with a unique correct interpretation ($\mathcal{U}$) and items with multiple correct interpretations ($\mathcal{M}$). But while the strict F1 score defined in Section 2.5 is appropriate for items in $\mathcal{U}$, it unfairly penalizes Builders on items in $\mathcal{M}$ when their structure is correct, but placed in a different location and/or orientation than the reference structure. A fair evaluation metric needs to account for these allowable changes in location and orientation for items in $\mathcal{M}$, i.e. structures placed on an empty board when the context does not uniquely identify their placement.
@@
To evaluate a model’s net action set $A_{M}^{\mathrm{net}}$ against a human reference net action set $A_{H}^{\mathrm{net}}$ for an item in $\mathcal{M}$, where $A_{M}$ leads from the (empty) board to a predicted structure $S_{M}$, and $A_{H}$ leads from the (empty) board to a reference structure $S_{H}$, we therefore first identify an optimal alignment of the predicted structure to the reference, $\mathcal{A}^{*}=\operatorname*{arg\,min}_{\mathcal{A}}(\Delta(\mathcal{A}(S_{M}),S_{H}))$, and apply the same transformation $\mathcal{A}^{*}$ to all actions in the model’s net action set $A_{M}^{\mathrm{net}}$, yielding an aligned action sequence $\widetilde{A}_{M}^{\mathrm{net}}$ that can now be fairly compared against $A_{H}^{\mathrm{net}}$ by our original strict F1 metric. That is, a fairer F1 metric computes a strict F1 score for the predicted action sequences $A_{M}^{\mathrm{net}}$ for any item in $\mathcal{U}$, because the reference sequences for these items yield the only correct structures for these items, but for any item in $\mathcal{M}$, it computes a strict F1 score for the optimally aligned predicted action sequences $\widetilde{A}_{M}^{\mathrm{net}}$ because this item has multiple correct interpretations that only differ from each other in location and orientation. Thus, in the example in Figure 4 (bottom), any predicted orange 3x3 U-shaped structure placed on the ground will receive a fairer F1 score of 1.0.
@@ 4.2 Auxiliary metrics
$\textbf{Location equivalence:}~(t,c,\langle x,y,z\rangle)\equiv_{\textrm{Loc}}(t^{\prime},c^{\prime},\langle x^{\prime},y^{\prime},z^{\prime}\rangle)$ $\iff\langle x,y,z\rangle=\langle x^{\prime},y^{\prime},z^{\prime}\rangle$
@@
$M^{\textrm{Type}}$ $=\{t\mid(t,c,\langle x,y,z\rangle)\in A^{\textrm{net}}\}$
$M^{\textrm{Color}}$ $=\{(t,c)\mid(t,c,\langle x,y,z\rangle)\in A^{\textrm{net}}\}$
$M^{\textrm{Loc}}$ $=\{\langle x,y,z\rangle\mid(t,c,\langle x,y,z\rangle)\in A^{\textrm{net}}\}$
which yields (strict) Type, Color and Location precision, recall and F1 scores by comparing the sizes of the corresponding multisets and their intersections (we only show Type F1 below, as Color and Location F1 are similarly defined):
$\textbf{Strict Type Precision}~P_{\textrm{Type}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{|(M_{M}^{\textrm{Type}}\cap M_{H}^{\textrm{Type}})|}{|M_{M}^{\textrm{Type}}|}$
$\textbf{Strict Type Recall}~R_{\textrm{Type}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{|(M_{M}^{\textrm{Type}}\cap M_{H}^{\textrm{Type}})|}{|M_{H}^{\textrm{Type}}|}$
$\textbf{Strict Type F1}~F1_{\textrm{Type}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{2\cdot P_{\textrm{Type}}\cdot R_{\textrm{Type}}}{P_{\textrm{Type}}+R_{\textrm{Type}}}$
Fairer counterparts to these strict metrics are defined analogously to the overall metric, and we will use the fair variants in the remainder of this paper. Intuitively, the Type metric captures how good a model is at understanding what type of actions to perform – placements, or removals, or both, and how many of each type. The Color metric is stricter than he Type metric, in that it quantifies how good a model is at understanding the colors of blocks that need to be placed or removed. The Location metric quantifies how good a model is at understanding the specific locations of blocks that need to be placed or removed (without regard to color), i.e., it helps to evaluate spatial reasoning.
@@
To evaluate how good a Builder model is at understanding the structure in which the required blocks need to be placed or removed, we define a metric, Shape F1, that is agnostic to the absolute $\langle x,y,z\rangle$ locations of the required blocks, but only cares that they be placed/removed in the right locations relative to each other. Essentially, this metric compares net actions modulo their exact placement and orientation. Similar to the methodology used for fairer F1 (Section 4.1), where we align the resulting built structures, we optimally align the model’s net actions $A_{M}^{\mathrm{net}}$ against the human reference net actions $A_{H}^{\mathrm{net}}$, yielding aligned net actions $\hat{A}_{M}^{\mathrm{net}}$ that can now be compared against $A_{H}^{\mathrm{net}}$ by our original strict F1 metric. This is done for all items in the v2 test set (and not just the ones with multiple interpretations like was done for fairer F1). We will see examples of Shape F1 and the other auxiliary metrics in Sections 7.2.
@@ 4.3 BAP v2 evaluation benchmark
The above metrics enable a fairer, robust, and more insightful evaluation. We validate this on the GRU-based baseline model for the BAP task (Section 2.6). Table 1 reports the full battery of scores included in the BAP v2 evaluation benchmark, i.e. overall and auxiliary F1 scores on the overall test set as well as for the empty and non-empty board (EB and NEB) subsets. We note that the overall F1 score has risen from 21.1 reported in Jayannavar, Narayan-Chen, and Hockenmaier (2020) to 27.3 under our newer, fairer evaluation, or 53.2 for EB and 24.2 on the (more frequent) NEB items. The distinction between EB vs. NEB is important since EB examples are generally easier, making EB performance a basic competency or sanity test, and a reflection of a model’s maximum potential quality. EB performance is also an indication of a model’s ability of a model to begin a game accurately – an elementary yet important trait to measure, especially for interactive/online settings where errors might cascade throughout a game. The auxiliary metrics provide further insight: in particular, we see that the model achieves reasonably good F1 scores of 65.2% for Type and 62.4% for Color, but only 27.7% for Location F1, which is also very close to the overall F1 of 27.3% (similar results hold both for the EB and NEB setting). This suggests that spatial reasoning is the key bottleneck for high performance on this task.
@@
Table 1: F1 scores for the GRU baseline model on the v2 benchmark
5 Revisiting BAP Training Data: Generating Synthetic Data
Synthetic Target Structures
Random Shape-Based
Refer to caption Refer to caption Refer to caption
Figure 8: Synthetic target structures
Synthetic BAP items
(A) Block-based dialogues for random targets ($D_{\mathrm{r}}$)
(B) Block-based dialogues for shape-based targets ($D_{\mathrm{sb}}$)
(C) Shape-based dialogues for shape-based targets ($D_{\mathrm{ss}}$)
Figure 9: We generate three different types of synthetic BAP items, resulting in datasets $D_{\mathrm{r}}$, $D_{\mathrm{sb}}$ and $D_{\mathrm{ss}}$
Our process to generate synthetic BAP items
The synthetic BAP item generated by the process above
Figure 10: Our process to generate synthetic BAP items (top) and a BAP item generated by that process (bottom)
Feature MDC ($D_{\mathrm{mc}}$) $D_{\mathrm{r}}$ $D_{\mathrm{sb}}$ $D_{\mathrm{ss}}$
Target Structure Human-designed, complex and meaningful (e.g., furniture, animals). Dynamically and randomly generated; typically disordered. Predefined composites of 6 elementary shapes. Predefined composites of 3 simpler shapes.
Architect Planning Complex, strategic & adaptive human planning with hierarchical decomposition of target structures based on goals of efficiency, naturalness in instruction giving, etc. Minimalistic; plans single random block placement/removal per turn. Mid-level; plans one shape at a time, decomposed into block-level instructions. High-level; plans construction of one full shape instance per turn.
Instruction Abstraction Multi-level: from high-level concepts (“flower") to low-level commands. Low-level; single block references only. Low-level; but more abstract than $D_{\mathrm{r}}$; decomposes shapes into block-level actions. Mid-level; refers to entire shapes (“row", “plane").
Semantic Naming Frequent use of names for commonplace objects (“chicken"), shapes & user-defined concepts (“r-windows"). N/A (block-level instructions only). N/A (block-level instructions only). Limited; template-based names for shapes (“pillar", “line").
Spatial Relations Highly diverse, complex, and nuanced (“parallel", “symmetrical", “hanging"). Simple 1D/2D relations anchored to arbitrary reference blocks. Simple 1D/2D relations anchored to the “last placed block". All 1D/2D/3D relations to specify shape’s starting point and Builder-anchored relations (“going to the left of you") to specify orientation.
Floating Blocks Present; instructions range from single-shot commands to step-by-step guidance. Present; single-shot commands only. Present; single-shot commands only. Present; single-shot commands only.
Context Clarity Variable; includes unclear/ambiguous contexts (Section 3). Always clear & unambiguous. Always clear & unambiguous. Always clear & unambiguous.
Dialogue Synchronicity Asynchronous; mutual interruptions by A and B are possible. Synchronous; discrete, structured turns. Synchronous; discrete, structured turns. Synchronous; discrete, structured turns.
Structural Correctness Contains correct & incorrect structures (due to mistakes/interruptions; Section 3). Always correct. Always correct. Always correct.
Dialogue Act Diversity High; wide mix (instructions, confirmations, corrections, social chat, etc.). Low; dominated by instructions & short clarification exchanges. Low; dominated by instructions & short clarification exchanges. Low; dominated by instructions and short clarification exchanges.
Builder Initiative High; B actively verifies, suggests, gives updates and extrapolates. Low; limited to simple clarification questions. Low; limited to simple clarification questions. Low; limited to simple clarification questions.
Clarification Exchanges Varied and natural questions about ambiguities ( “is it flat?"). Simulated; on missing block color or location. Simulated; on missing block color or location. Simulated; on missing color, location, size, or orientation.
Referring Expressions Highly varied; includes implicit arguments and ellipsis (“two more in the same direction"). Simple; disambiguate blocks by color and relative position (“leftmost red block"). Simple; refers to the last placed block as spatial anchor. Simple; refers to last placed block of previous shape as spatial anchor.
Builder’s Low-Level Planning Required & involved; scales with instruction abstraction. Minimal; executes single-block instructions. Minimal; executes block-level instructions. Required but simple; uses heuristics (e.g., column-by-column, row-by-row, or zigzag patterns) for shape instructions.
Table 2: Detailed qualitative comparison of MDC and synthetic datasets. For all rows, MDC features are a superset of the ones for the synthetic datasets.
Despite the data augmentation (Section 2.3), the BAP task suffers from relatively limited training data, but collecting a significantly larger amount of human data would have been prohibitively expensive and impractical. More importantly, while the MDC dialogues are far more linguistically rich and varied than our synthetic data, the baseline models struggle even with basic spatial understanding and reasoning (as reflected in the Shape and Location F1 scores). In order to achieve comprehensive coverage of various spatial relations, we therefore opted to create synthetic training data that would allow us to train our models in these basic skills. Specifically, we distinguish between two different types of target structures ––random and shape-based (Figure 8) –, and three different types of dialogues ––block-based for random or shape-based targets, or shape-based for shape-based targets–– (Figure 9), described in detail below.
@@ 5 Revisiting BAP Training Data: Generating Synthetic Data
We design novel Minecraft target structure and dialogue simulators that emulate the MCBT, and use them to generate a set of three synthetic dialogue datasets with corresponding target structures. Each dataset is modeled after the MDC and, while understandably simpler, preserves essential elements of realistic human behavior observed in the MDC as much as possible (including, e.g. clarification exchanges, Architect A ’s and Builder B ’s planning perspectives, dialogue flow, etc.). While we specifically emphasize spatial relations (as it is a core motivation) and representing examples of different complexities, we also vary the level of instruction abstraction, the complexity of referring expressions, the complexity of target structures, etc. across these datasets. Each is thus unique, contributing to the diversity and richness of the overall set.
@@
Looking ahead, such data could also be crucial for training more sophisticated, data-hungry deep transformer models and training/fine-tuning increasingly large LLMs. And although our primary focus is on BAP, the simulators and data introduced here are applicable to other MCBT subtasks as well (e.g. the AUG subtask; Section 2.1.2), supporting advancements across the broader MCBT ecosystem. We discuss this in more detail in Section 8.2.
5.1 Our dialogue simulation framework
As described in Section 2.1, the MCBT is characterized by asynchronous, loosely structured dialogue between human players with minimal constraints on their language and actions. To make the simulation of this complex interaction tractable and to facilitate a simplified data generation process, we introduce a simpler and more structured framework. Like the original MDC data, for each game, a target structure and a complete log that interleaves the dialogue between A and B (consisting of instructions, potential clarification exchanges), B’s actions (which reflect the evolving structure being built), and B’s changing position and orientation (so that spatial references in the utterances can be correctly interpreted). Our simulators generate grounded synthetic dialogues that meet these desiderata by iteratively simulating the MCBT until the target structure is built. Each iteration of the algorithm produces a BAP item, $(\mathcal{H}^{t-1},S^{t-1},B^{t},D^{t},A^{t},S^{t})$ (as defined in Section 2.2), by simulating the dialogue $D^{t}$ and actions $A^{t}$ following $S^{t-1}$ to produce the next structure, $S^{t}$. Each item consists of the specification of B’s current position $B^{t}$, a non-empty dialogue segment $D^{t}$A’s instruction, optional clarification exchanges with B, and confirmations– followed by B’s actions $A^{t}$ that alter the built structure from $S^{t-1}$ to $S^{t}$, thus constructing part of the target structure. $\mathcal{H}^{t-1}$ and $S^{t-1}$ represent the current game context that is input to this iteration of the algorithm.
The target structure for the dialogue is handled in one of two ways: it is either provided as a predefined input to the algorithm (generated by a separate target structure simulator) or generated dynamically. If predefined, each iteration involves selecting the next blocks to place from this target. If generated dynamically, each iteration involves choosing blocks randomly to place or remove, and the final constructed configuration is designated as the target structure for the dialogue post-hoc.
Each iteration proceeds according to the steps detailed below, which is also visualized in the running example in Figure 10 (bottom).